Use the following linear regression equation to answer the questions.

x1 = 1.4 + 3.7x2 – 8.3x3 + 1.8x4

(c) If x2 = 2, x3 = 6, and x4 = 10, what is the predicted value for x1? (Use 1 decimal place.)

Suppose x3 and x4 were held at fixed but arbitrary values and x2 increased by 1 unit. What would be the corresponding change in x1?


Suppose x2 increased by 2 units. What would be the expected change in x1?


Suppose x2 decreased by 4 units. What would be the expected change in x1?


(e) Suppose that n = 16 data points were used to construct the given regression equation and that the standard error for the coefficient of x2 is 0.384. Construct a 99% confidence interval for the coefficient of x2. (Use 2 decimal places.)

lower limit
upper limit


(f) Using the information of part (e) and level of significance 1%, test the claim that the coefficient of x2 is different from zero. (Use 2 decimal places.)

t
t critical ±
Answer #1

c)

predicted value for x1 =1.4+3.7*2-8.3*6+1.8*10=-23

for  x2 increased by 1 unit; corresponding change in x1 =3.7

for  x2 increased by 2 unit; corresponding change in x1 =7.4

for  x2 increased by 4 unit; corresponding change in x1 =14.8

e)

for 99% CI ; critical t=3.055

lower limit =estimate -t*std error =3.7-0.384*3.055=2.53

upper limit=estimate +t*std error =3.7+0.384*3.055=4.87

f)

t=estimate/std error =3.7/0.384=9.64

tcitical =-/+ 3.055 ~3.05